OpenStudy (anonymous):

A circle is centered at (7, 8) and has a radius of 11. Which of the following is the equation for this circle? (x − 7)^2 + (y − 8)^2 = 121 (x − 7)^2 + (y − 8)^2 = 11 (x + 7)^2 + (y + 8)^2 = 121 (x + 7)^2 + (y + 8)^2 = 11

OpenStudy (anonymous):

will donate medal to best answer

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

@jim_thompson5910 @mathstudent55 @ganeshie8 @misty1212

OpenStudy (anonymous):

hi :)

OpenStudy (misty1212):

make a guess, i bet you are right general form is \[(x-h)^2+(y-k)^2=r^2\] put \(h=7,k=8,r=11\)

OpenStudy (anonymous):

so (x-7)^2+(y-8)^2=11^2 ?

OpenStudy (misty1212):

yeah although they need you to square 11

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so (x^2+49)+(y+64)=121 ?

OpenStudy (anonymous):

@misty1212 ?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

D:

OpenStudy (anonymous):

oh no!!

OpenStudy (anonymous):

can you help me ?

OpenStudy (anonymous):

just take what you wrote above

OpenStudy (anonymous):

\[(x-7)^2+(y-8)^2=11^2 \]

OpenStudy (anonymous):

(x^2+49)+(y+64)=121 ?

OpenStudy (anonymous):

all @misty1212 was saying is that in your answer choices the 11 is squared

OpenStudy (anonymous):

so does that matter or not ?

OpenStudy (freckles):

are you trying to expand the square parts?

OpenStudy (anonymous):

no it doesn't matter except for your answer choices

OpenStudy (anonymous):

i dont know im bad at geometry

OpenStudy (freckles):

\[(x-7)^2=(x-7)(x-7)=x(x-7)-7(x-7)=x^2-7x-7x+49 \\ =x^2-14x+49\] no your choices do not have this form

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

D: D: D:

OpenStudy (freckles):

\[(x-h)^2+(y-k)^2=r^2 \\ \text{ you are give} (h,k)=(7,8) \text{ and } r=11 \\ \text{ just plug into this form }\] I think @satellite73 already walked you though this

OpenStudy (freckles):

replace h with 7 replace k with 8 replace r with 11

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(x-7)^2+(y-8)^2=121 ?

OpenStudy (freckles):

that is right

OpenStudy (anonymous):

tnx u deserve the medal